# LitmusKt: Concurrency Stress Testing for Kotlin

### v1

@@ Abstract.
We present LitmusKt—the first tool for litmus testing concurrent programs in Kotlin. The tool’s novelty also lies in the fact that Kotlin is a multiplatform language, i.e., it compiles into multiple platforms, which means that the concurrency has to be tested on several of them. Our tool allows writing litmus tests in a single custom DSL, and these tests are then run in Kotlin/Native and Kotlin/JVM, two main platforms for concurrent programming in Kotlin. Using LitmusKt, we discovered novel bugs in the Kotlin compiler, which we then fixed and they are no longer present. Moreover, LitmusKt was integrated into the CI pipeline for Kotlin. We believe that our tool is valuable for further studying concurrency in Kotlin and other multiplatform languages, as well as for further developing the Kotlin memory model.
LitmusKt is openly available on GitHub: https://github.com/Jetbrains-Research/litmuskt. The demonstration video is available on YouTube: https://youtu.be/gXI0aYJDnRw.
@@ 1. Introduction
As a classic example, consider the Store Buffering program in Figure 1, where the program first initializes locations $x$, $y$, $r1$, and $r2$ to 00 and then spawns two threads. Suppose we try to run this program on a single-core machine without compiler optimizations. This would result in some interleaving of read and write operations, as defined by the model of sequential consistency (SC) (Lamport, 1979). Notice that both threads start with a write operation either to location $x$ or to location $y$, implying that $r1$ and $r2$ both being equal to 00 at the end of execution is seemingly impossible. However, such behavior can be observed in a real-world multi-core setting due to compilers reordering instructions in threads (e.g., GCC may do so (Vafeiadis et al., 2015)), and such outcomes are referred to as weak (Alglave et al., 2014).
@@
LitmusKt is openly available on GitHub: https://github.com/Jetbrains-Research/litmuskt. The demonstration video is available on YouTube: https://youtu.be/gXI0aYJDnRw.
@@ 2.1. Example of a LitmusKt Test
  • acceptable—sequentially consistent (SC) outcomes, i.e., reproducible as an interleaving of threads;
@@ 2.3. Kotlin/JVM Test Runner Based on jcstress
The ability of running tests on both Kotlin/Native and Kotlin/JVM and comparing their results was one of the main requirements of LitmusKt. At the same time, for Kotlin/JVM compiler, reusing the established tool jcstress (OpenJDK, 2024) would be advantageous, as it already implements many advanced techniques and has a proven track record of uncovering concurrency bugs in various JVM implementations (Aleksey Shipilev, 2024). Therefore, we decided to support jcstress in the LitmusKt as a backend for running tests on Kotlin/JVM.
@@
As demonstrated, LitmusKt conforms to all the requirements that we set forth. It allows writing litmus tests in a flexible yet concise syntax, runs them using efficient techniques, and supports multiple platforms by delegating to jcstress on JVM and using a custom runner for Native, unifying them under a common interface. Let us now highlight some discoveries we made with the help of LitmusKt.
@@ 3. Analysis
UPUB ACC FORB FORB FORB
@@
We ran a number of relatively standard (Moiseenko et al., 2021) litmus tests with LitmusKt on two machines: x86 machine (Intel i7-12700K CPU) and Arm machine (Apple M1 CPU). The results are presented in Table 2. Many of the tests that theoretically could show some weak behavior did not do so in practice. This is expected, because Kotlin memory model is not formalized, and as such, it cannot be determined in advance which tests may show weak behaviors in practice. However, some tests still showed noteworthy behaviors, and we will now discuss those tests in more detail.
Firstly, the Unsafe Publication test (UPUB for short) demonstrates forbidden outcomes. Its schematic description is shown in Figure 3. Here, one thread initializes a class instance, while another thread simultaneously tries to read its field. When we run this test, we expect to see either no instance in the reading thread (i.e., null), or a fully initialized instance, or a partially initialized class with default values for fields. However, as shown in Table 3, when running with Kotlin/Native 1.9.0 and on an Arm processor, it is possible to see ”garbage” values. The problem is exacerbated if we substitute the Int field with a class reference, because when we try to read a reference with a ”garbage” address, we get a segmentation fault.
@@
Moreover, the story with UPUB does not end there. We found more problems regarding this test, caused by the fact that classes can be published with fields equal to default values. For example, Kotlin has nullable and non-nullable types, and under normal conditions, it should not be possible to read null from a non-nullably-typed variable. However, a similar UPUB with reference test can be used to observe the opposite, and we do in fact see it. This fact presents a new challenge for language design.
@@ 4. Future work
Another scope for future work is adding further test runner optimizations, most notably, preventing false sharing. The term false sharing refers to a situation when independent variables get allocated on a single CPU cache line, which can easily prevent many weak behaviors. In our internal experiments, we have seen that using some method to avoid false sharing can significantly boost the frequency of weak behaviors, but unfortunately, we struggled to find a general solution to add into LitmusKt. We proposed adding a Kotlin/Native annotation similar to JVM’s @Contended, but the idea did not get much attention (Moiseenko, 2024). Still, this unsolved problem leaves room for LitmusKt improvement.

### v2

@@ Abstract.
We present LitmusKt—the first tool for litmus testing concurrent programs in Kotlin. The tool’s novelty also lies in the fact that Kotlin is a multiplatform language, i.e., it compiles into multiple platforms, which means that the concurrency has to be tested on several of them. Our tool allows writing litmus tests in a single custom DSL, and these tests are then run in Kotlin/Native and Kotlin/JVM, two main platforms for concurrent programming in Kotlin. Using LitmusKt, we discovered novel bugs in the Kotlin compiler, which we then fixed and they are no longer present. Moreover, LitmusKt was integrated into the CI pipeline for Kotlin. LitmusKt is available on GitHub: https://github.com/JetBrains-Research/litmuskt. The demo is available on YouTube: https://youtu.be/oWCZp_Huwss.
Kotlin, Litmus, Stress Testing, Concurrency, Compiler Design
journalyear: 2025copyright: acmlicensedconference: 33rd ACM International Conference on the Foundations of Software Engineering; June 23–28, 2025; Trondheim, Norwaybooktitle: 33rd ACM International Conference on the Foundations of Software Engineering (FSE Companion ’25), June 23–28, 2025, Trondheim, Norwaydoi: 10.1145/3696630.3728584isbn: 979-8-4007-1276-0/2025/06ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Concurrent programming structures
@@ 1. Introduction
As a classic example, consider the Store Buffering program in Figure 1, where the program first initializes locations $x$, $y$, $r1$, and $r2$ to 00 and then spawns two threads. Suppose we try to run this program on a single-core machine without compiler optimizations. This would result in some interleaving of read and write operations, as defined by the model of sequential consistency (Lamport, 1979). Notice that both threads start with a write operation either to location $x$ or to location $y$, implying that $r1$ and $r2$ both being equal to 00 at the end of execution is seemingly impossible. However, such behavior can be observed in a real-world multi-core setting due to compilers reordering instructions in threads (e.g., GCC may do so (Vafeiadis et al., 2015)), and such outcomes are referred to as weak (Alglave et al., 2014).
@@
LitmusKt is openly available on GitHub: https://github.com/JetBrains-Research/litmuskt. The demonstration video is available on YouTube: https://youtu.be/oWCZp_Huwss.
@@ 2.1. Example of a LitmusKt Test
  • acceptable—sequentially consistent outcomes, i.e., reproducible as an interleaving of threads;
@@ 2.3. Kotlin/JVM Test Runner Based on jcstress
The ability of running tests on both Kotlin/Native and Kotlin/JVM and comparing their results was one of the main requirements of LitmusKt. At the same time, for Kotlin/JVM compiler, reusing the established tool jcstress (OpenJDK, 2024) would be advantageous, as it already implements many advanced techniques and has a proven track record of uncovering concurrency bugs in various JVM implementations (Shipilev, 2024). Therefore, we decided to support jcstress in the LitmusKt as a backend for running tests on Kotlin/JVM.
@@
As demonstrated, LitmusKt conforms to all the requirements that we set forth. It allows writing litmus tests in a flexible yet concise syntax, runs them using efficient techniques, and supports multiple platforms by delegating to jcstress on JVM and using a custom runner for Native, unifying them under a common interface.
@@ 3. Analysis
UPUB ACC FORB ACC ACC
@@
We ran a number of relatively standard litmus tests (Moiseenko et al., 2021) with LitmusKt on two machines: x86 machine (Intel i7-12700K CPU) and Arm machine (Apple M1 CPU). The results are presented in Table 2. Many of the tests that theoretically could show some weak behavior did not do so in practice. This is expected, because Kotlin memory model is not formalized, and as such, it cannot be determined in advance which tests may show weak behaviors in practice. However, some tests still showed noteworthy behaviors, and we will now discuss those tests in more detail.
Firstly, the Unsafe Publication test (UPUB for short) demonstrates forbidden outcomes. Its schematic description is shown in Figure 3. Here, one thread initializes a class instance, while another thread simultaneously tries to read its field. When we run this test, we expect to see either no instance in the reading thread (i.e., null), or a fully initialized instance, or a partially initialized class with default values for fields. However, as shown in Table 3, when running with Kotlin/Native 1.9.0 on an Arm processor, it is possible to see ”garbage” values. The problem is exacerbated if we substitute the Int field with a class reference, because when we try to read a reference with a ”garbage” address, we get a segmentation fault.
@@
We found more problems regarding UPUB, caused by the fact that classes can be published with fields equal to default values. For example, Kotlin has nullable and non-nullable types, and under normal conditions, it should not be possible to read null from a non-nullably-typed variable. However, a similar UPUB with reference test can be used to observe the opposite, and we do in fact see it. This fact presents a new challenge for language design.
@@ 4. Future work
Another scope for future work is adding further test runner optimizations, most notably, preventing false sharing. The term false sharing refers to a situation when independent variables get allocated on a single CPU cache line, which can easily prevent many weak behaviors. In our internal experiments, we have seen that using some method to avoid false sharing can significantly boost the frequency of weak behaviors, but unfortunately, we struggled to find a general solution to add into LitmusKt. We proposed adding a Kotlin/Native annotation similar to JVM’s @Contended (Moiseenko, 2024).